# Sayantan Roy Sankha S. Basu Mihir K. Chakraborty School of Cognitive Science

### v1

@@ 1 Introduction
A logical structure (see [3, p. 84]) is a pair $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$, where $\mathscr{L}$ is a set, $\mathcal{P}(\mathscr{L})$ denotes the power set of $\mathscr{L}$ and |-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,\subseteq\,\mathcal{P}(\mathscr{L})\times\mathscr{L}$. Alternatively, a logical structure may be defined as a pair $(\mathscr{L},W)$, where $\mathscr{L}$ is a set and $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ is called a consequence operator. These two definitions are equivalent in the following sense.
@@ 2. $\bullet$
    Every logical structure of the form $(\mathscr{L},W)$ induces a logical structure of the form (,|-$(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{W})$ where |-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{W}\,\subseteq\,\mathcal{P}(\mathscr{L})\times\mathscr{L}$ and for all $\alpha\in\mathscr{L}$
@@
Suszko’s Thesis ($\mathbf{ST}$), named after the Polish logician Roman Suszko, is the (philosophical) claim that “there are but two logical values, true and false” (see [11, p. 169]). This thesis has been given formal contents by the so-called Suszko Reduction (the particular case of Łukasiewicz’s 3-valued logic Ł$\textbf{\L}_{3}$ is outlined in [37, pp. 378--379] by Suszko himself). However, as has been pointed out in [10, p. 6], “there seems to be no paper where Suszko explicitly formulates (SR)[the Suszko Reduction] in full generality!” Furthermore, there seems to be no explicit definition of ‘logical value’. This makes $\mathbf{ST}$ rather ambiguous, difficult to justify and open to interpretation. The situation is explained succinctly in [7, p. 293] as follows (emphasis ours).
@@
Interpretations differ. For example, the ‘real’ reason behind the existence of adequate bivalent semantics for every Tarski-type logical structure -- in other words, the ‘logical 2-valuedness’ of Tarski-type logical structures -- has been explained by Malinowski in [27, p. 80] as follows.
@@
Thus, according to him, ‘logical value’ corresponds to the number of divisions of the universe of interpretation (cf. [38] and [40]). This particular conceptualisation of ‘logical value’ is termed by Malinowski as inferential value. This notion of many-valuedness is different from the more traditional conception of many-valuedness -- algebraic many-valuedness [^2] Sometimes ‘algebraic many-valuedness’ is also referred to as ‘referential many-valuedness’ (see, e.g., [28, 29]). However, we will not use this terminology in this chapter.. The difference between these is explained in [37, p. 378] as follows (emphasis ours).
@@ 2 Fundamental Notions
Suszko avoids speaking of truth values. Instead, one of his main points is to distinguish between algebraic values (the members of the algebra of truth values, introduced by Łukasiewicz and others and usually interpreted as such in the literature) and logical values (of which his intuition admits only two, true and false). The mathematical result $\ldots$ can be paraphrased $\ldots$ only after the adoption of a peculiar conception of what is a “two-valued semantics”, and in particular of what is, or can be, a “semantics”$\ldots$
@@
If, in particular, $I=\{1,\ldots,n\}$ with $n\in\mathbf{N}\setminus\{0\}$, instead of writing $(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ we write $(\mathbf{M},\models_{1},\ldots,\models_{n},S,\mathcal{P}(\mathscr{L}))$. If $(m,\Gamma)\in\,\models_{i}$, we write $m\models_{i}\Gamma$. Given a semantics $\mathfrak{S}=(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ the entailment relation induced by $\mathfrak{S}$, |-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}\,\subseteq\mathcal{P}(\mathscr{L})\times\mathscr{L}$ is defined as follows.
@@
A logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ (respectively, $(\mathscr{L},W)$) is said to be adequate with respect to a semantics $\mathfrak{S}$ if |-=|-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$ (respectively, $W=W_{\mathfrak{S}}$; where $W_{\mathfrak{S}}$ is the consequence operator induced by |-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$). A logical structure is said to have an adequate semantics if there exists a semantics $\mathfrak{S}$ for $\mathscr{L}$ such that |-=|-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$ (respectively, $W=W_{\mathfrak{S}}$).
@@
Then, one can show that (,|-$(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{B}_{\mathcal{V}}})$ is of Tarski-type. In fact, given a Tarski-type logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$, one can prove that |-=|-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{B}_{\mathcal{S}}}$, where $\mathcal{S}=\{\chi_{W^{\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}}(\Sigma)}\in\{0,1\}^{\mathscr{L}}:\Sigma\subseteq\mathscr{L}\}$.
@@
Then, one can easily show that (,|-$(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{P}_{\mathcal{V}}})$ is of Tarski-type. In fact, given a Tarski-type logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$, one can prove that |-=|-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{P}_{\mathcal{S}}}$ where $P=\{0,1\},\preceq\,=\{(0,0),(0,1),(1,1)\}$, and $\mathcal{S}$ is as defined in the previous example.
@@
Given a set $\mathscr{L}$, and a semantics $\mathfrak{S}$ for $\mathscr{L}$, the entailment relation |-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$ as we have defined above, is not the only possible one. Below we mention two other possibilities.
@@
Generally, the notion of semantics is not defined explicitly. A notable exception [^4] Equivalent notions under the names of model-based structure and abstract model logic have been defined in [39, p. 106] and [21, p. 21] respectively. is [4, pp. 107--108] where a semantics is defined as a triple $\langle\mathfrak{F};\boldsymbol{U};\text{mod}\rangle$, where
@@
    |-=|-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$.
@@
    For all algebraically $\mu$-valued semantics $\mathfrak{T}=(\mathbf{M}^{\prime},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ with $\mu<\kappa$, |-|-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,\neq\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{T}}$.
@@
    |-=|-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$.
@@
    For all functionally $\mu$-valued semantics $\mathfrak{T}=(\mathbf{M}^{\prime},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ with $\mu<\kappa$, |-|-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,\neq\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{T}}$.
@@ 3 Monotonic Logical Structures: A Case Study
Let $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S,\mathcal{P}(\mathscr{L}))$, where $S=\{(\models_{1},\models_{2})\}$, be a strongly granular semantics for $\mathscr{L}$. Then $W_{\mathfrak{S}}$ is monotonic.
@@
If possible, suppose $\Gamma\cup\Sigma\subseteq\mathscr{L}$ be such that $\Gamma\subseteq\Sigma$ but $W_{\mathfrak{S}}(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Sigma)$. Hence, there exists $\alpha\in W_{\mathfrak{S}}(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Sigma)$. Since $\alpha\notin W_{\mathfrak{S}}(\Sigma)$, there exists $m\in\mathbf{M}$ such that $m\models_{1}\Sigma$ but $m\not\models_{2}\{\alpha\}$. However, since $\Gamma\subseteq\Sigma$, $m\models_{1}\Sigma$ and $\mathfrak{S}$ is strongly granular -- and hence granular at the 1st component -- it follows that $m\models_{1}\Gamma$ as well. But then, since $\alpha\in W_{\mathfrak{S}}(\Gamma)$, we must have $m\models_{2}\{\alpha\}$- a contradiction. Thus, $W_{\mathfrak{S}}(\Gamma)\subseteq W_{\mathfrak{S}}(\Sigma)$. Hence, $W_{\mathfrak{S}}$ is monotonic. ∎
@@ 3. $\bullet$
If $W(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Gamma)$, then there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. So, there exists $\mu_{\Sigma}\in\mathbf{M}$ such that $\mu_{\Sigma}\models_{1}\Gamma$ but $\mu_{\Sigma}\not\models_{2}\{\alpha\}$. Now, $\mu_{\Sigma}\models_{1}\Gamma$ implies that $\mu_{\Sigma}(\Gamma)\subseteq\{1,2\}$, or, equivalently $\Gamma\subseteq\Sigma$. Then, since $W$ is monotonic, we have $W(\Gamma)\subseteq W(\Sigma)$. Consequently, $\alpha\in W(\Gamma)$ implies that $\alpha\in W(\Sigma)$. But, $\alpha\in W(\Sigma)$ is equivalent to requiring that $\mu_{\Sigma}(\alpha)\in\{1,3\}$, i.e., $\mu_{\Sigma}\not\models_{2}\{\alpha\}$- a contradiction. Consequently, $W(\Gamma)\subseteq W_{\mathfrak{S}}(\Gamma)$.
@@ 3.1 $q$-consequence Operators
    $W^{\infty}(\Gamma)\subseteq W(\Gamma)\cup\Gamma$br2sinceWsatisfies quasi-closure
    $\Gamma\subseteq W(\Gamma)\cup\Gamma\subseteq W^{\infty}(\Gamma)$br2by(2)
@@
    $\underline{(3)\implies(2):}$ Suppose that $W$ is not monotonic. So, there exists $\Gamma\cup\Sigma\subseteq\mathscr{L}$ with $\Gamma\subseteq\Sigma$ such that $W(\Gamma)\not\subseteq W(\Sigma)$. This implies that there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W(\Sigma)$. Since $\Gamma\subseteq\Sigma\subseteq W^{\infty}(\Sigma)$ and $\Sigma$ itself is downward $q$-closed by (3), it follows that $W(\Gamma)\subseteq W(\Sigma)$. Hence, $\alpha\in W(\Gamma)$ implies that $\alpha\in W(\Sigma)$ as well -- a contradiction. Therefore, $W$ is monotonic.
@@
    $\underline{(4)\implies(2):}$ We first show that $W$ is monotonic. For this, suppose that there exists $\Gamma\cup\Sigma\subseteq\mathscr{L}$ with $\Gamma\subseteq\Sigma$ such that $W(\Gamma)\not\subseteq W(\Sigma)$. This implies that there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W(\Sigma)$. Consequently, by (4), there exists $\Delta\supseteq\Sigma$ such that $\Delta$ is downward $q$-closed and $\alpha\notin W(\Delta)$. Now, since $\Gamma\subseteq\Sigma\subseteq\Delta\subseteq W^{\infty}(\Delta)$, and $\Delta$ is downward $q$-closed, $W(\Gamma)\subseteq W(\Delta)$. So, since $\alpha\in W(\Gamma)$, $\alpha\in W(\Delta)$ as well -- a contradiction. Hence, $W$ is monotonic.
    We now show that $W$ satisfies quasi-closure as well. Suppose the contrary, i.e., $W(W^{\infty}(\Gamma))\neq W(\Gamma)$. Then we must have $W\left(W^{\infty}(\Gamma)\right)\not\subseteq W(\Gamma)$ for some $\Gamma\subseteq\mathscr{L}$. So, there exists $\alpha\in W\left(W^{\infty}(\Gamma)\right)$ such that $\alpha\notin W(\Gamma)$. Then, by (4), there exists a downward $q$-closed $\Sigma\supseteq\Gamma$ such that $\alpha\notin W(\Sigma)$. Since $\Sigma$ is downward $q$-closed and $W^{\infty}(\Sigma)\subseteq W^{\infty}(\Sigma)$, it follows that $\alpha\notin W\left(W^{\infty}(\Sigma)\right)$. Now, as $W$ is monotonic, $\Gamma\subseteq\Sigma$ implies that $W^{\infty}(\Gamma)\subseteq W^{\infty}(\Sigma)$, and hence $W\left(W^{\infty}(\Gamma)\right)\subseteq W\left(W^{\infty}(\Sigma)\right)$. Thus, $\alpha\notin W\left(W^{\infty}(\Gamma)\right)$ as well -- a contradiction. Hence, $W\left(W^{\infty}(\Gamma)\right)\subseteq W(\Gamma)$. This proves (2).
@@
    The proof that every structural $q$-type logic has a 3-valued description is already been given, e.g., in [30, pp. 204--206]. However, a similar argument works for arbitrary $q$-type logical structures as well. Below we give a proof of the same, to keep the presentation self-contained.

### v2

# arXiv:2408.13769v2[math.LO] 01 Mar 2026

@@ 1 Introduction
A logical structure (see [3, p. 84]) is a pair $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$, where $\mathscr{L}$ is a set, $\mathcal{P}(\mathscr{L})$ denotes the power set of $\mathscr{L}$ and $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,\subseteq\,\mathcal{P}(\mathscr{L})\times\mathscr{L}$. Alternatively, a logical structure may be defined as a pair $(\mathscr{L},W)$, where $\mathscr{L}$ is a set and $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ is called a consequence operator. These two definitions are equivalent in the following sense.
@@ 2. $\bullet$
    Every logical structure of the form $(\mathscr{L},W)$ induces a logical structure of the form $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{W})$ where $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{W}\,\subseteq\,\mathcal{P}(\mathscr{L})\times\mathscr{L}$ and for all $\alpha\in\mathscr{L}$
@@
Suszko’s Thesis ($\mathbf{ST}$), named after the Polish logician Roman Suszko, is the (philosophical) claim that “there are but two logical values, true and false” (see [11, p. 169]). This thesis has been given formal contents by the so-called Suszko Reduction (the particular case of Łukasiewicz’s 3-valued logic $\textbf{Ł}_{3}$ is outlined in [37, pp. 378-379] by Suszko himself). However, as has been pointed out in [10, p. 6], “there seems to be no paper where Suszko explicitly formulates (SR)[the Suszko Reduction] in full generality!” Furthermore, there seems to be no explicit definition of ‘logical value’. This makes $\mathbf{ST}$ rather ambiguous, difficult to justify and open to interpretation. The situation is explained succinctly in [7, p. 293] as follows (emphasis ours).
@@
Interpretations differ. For example, the ‘real’ reason behind the existence of adequate bivalent semantics for every Tarski-type logical structure - in other words, the ‘logical 2-valuedness’ of Tarski-type logical structures - has been explained by Malinowski in [27, p. 80] as follows.
@@
Thus, according to him, ‘logical value’ corresponds to the number of divisions of the universe of interpretation (cf. [38] and [40]). This particular conceptualisation of ‘logical value’ is termed by Malinowski as inferential value. This notion of many-valuedness is different from the more traditional conception of many-valuedness - algebraic many-valuedness [^2] Sometimes ‘algebraic many-valuedness’ is also referred to as ‘referential many-valuedness’ (see, e.g., [28, 29]). However, we will not use this terminology in this chapter.. The difference between these is explained in [37, p. 378] as follows (emphasis ours).
@@ 2 Fundamental Notions
[…]
@@
If, in particular, $I=\{1,\ldots,n\}$ with $n\in\mathbf{N}\setminus\{0\}$, instead of writing $(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ we write $(\mathbf{M},\models_{1},\ldots,\models_{n},S,\mathcal{P}(\mathscr{L}))$. If $(m,\Gamma)\in\,\models_{i}$, we write $m\models_{i}\Gamma$. Given a semantics $\mathfrak{S}=(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ the entailment relation induced by $\mathfrak{S}$, $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}\,\subseteq\mathcal{P}(\mathscr{L})\times\mathscr{L}$ is defined as follows.
@@
[…]
@@
Then, one can show that $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{B}_{\mathcal{V}}})$ is of Tarski-type. In fact, given a Tarski-type logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$, one can prove that $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{B}_{\mathcal{S}}}$, where $\mathcal{S}=\{\chi_{W^{\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}}(\Sigma)}\in\{0,1\}^{\mathscr{L}}:\Sigma\subseteq\mathscr{L}\}$.
@@
Then, one can easily show that $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{P}_{\mathcal{V}}})$ is of Tarski-type. In fact, given a Tarski-type logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$, one can prove that $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{P}_{\mathcal{S}}}$ where $P=\{0,1\},\preceq\,=\{(0,0),(0,1),(1,1)\}$, and $\mathcal{S}$ is as defined in the previous example.
@@
Given a set $\mathscr{L}$, and a semantics $\mathfrak{S}$ for $\mathscr{L}$, the entailment relation $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$ as we have defined above, is not the only possible one. Below we mention two other possibilities.
@@
Generally, the notion of semantics is not defined explicitly. A notable exception [^4] Equivalent notions under the names of model-based structure and abstract model logic have been defined in [39, p. 106] and [21, p. 21] respectively. is [4, pp. 107-108] where a semantics is defined as a triple $\langle\mathfrak{F};\boldsymbol{U};\text{mod}\rangle$, where
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Notice that both Definition 2.9 and Definition 2.13 can be understood independent of the precise definition of $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$. Hence, even though, e.g., Definition 2.13 implies that the every inferentially $\kappa$-valued logical structure is necessarily monotonic  [^5] A special thanks to Laurent Dion for pointing this out. (see Theorem 3.1 for the proof) – hence, inferentially 4-valued, by Corollary 3.41 – it is possible to formulate other definitions of ‘inferentially $\kappa$-valued logical structure’ (depending on different definitions of $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$), that are not necessarily so. Similar remarks apply to Definition 2.9.
Remark 2.18.
@@ 3 Monotonic Logical Structures: A Case Study
Let $\mathfrak{S}=(\textbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$, be a strongly granular semantics for $\mathscr{L}$. Then $W_{\mathfrak{S}}$ is monotonic.
@@
If possible, suppose $\Gamma\cup\Sigma\subseteq\mathscr{L}$ be such that $\Gamma\subseteq\Sigma$ but $W_{\mathfrak{S}}(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Sigma)$. Hence, there exists $\alpha\in W_{\mathfrak{S}}(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Sigma)$. Since $\alpha\notin W_{\mathfrak{S}}(\Sigma)$, there exists $(i,j)\in I\times I$ with $(\models_{i},\models_{j})\in S$ and $m\in\mathbf{M}$ such that $m\models_{i}\Sigma$ but $m\not\models_{j}\{\alpha\}$. However, since $\Gamma\subseteq\Sigma$, $m\models_{i}\Sigma$ and $\mathfrak{S}$ is strongly granular, it follows that $m\models_{i}\Gamma$ as well. But then, since $\alpha\in W_{\mathfrak{S}}(\Gamma)$ and $(\models_{i},\models_{j})\in S$, we must have $m\models_{j}\{\alpha\}$ - a contradiction. Thus, $W_{\mathfrak{S}}(\Gamma)\subseteq W_{\mathfrak{S}}(\Sigma)$. Hence, $W_{\mathfrak{S}}$ is monotonic. ∎
@@ 3. $\bullet$
If $W(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Gamma)$, then there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. So, there exists $\mu_{\Sigma}\in\mathbf{M}$ such that $\mu_{\Sigma}\models_{1}\Gamma$ but $\mu_{\Sigma}\not\models_{2}\{\alpha\}$. Now, $\mu_{\Sigma}\models_{1}\Gamma$ implies that $\mu_{\Sigma}(\Gamma)\subseteq\{1,2\}$, or, equivalently $\Gamma\subseteq\Sigma$. Then, since $W$ is monotonic, we have $W(\Gamma)\subseteq W(\Sigma)$. Consequently, $\alpha\in W(\Gamma)$ implies that $\alpha\in W(\Sigma)$. But, $\alpha\in W(\Sigma)$ is equivalent to requiring that $\mu_{\Sigma}(\alpha)\in\{1,3\}$, i.e., $\mu_{\Sigma}\models_{2}\{\alpha\}$ - a contradiction. Consequently, $W(\Gamma)\subseteq W_{\mathfrak{S}}(\Gamma)$.
@@ 3.1 $q$-consequence Operators
    $\subseteq W(W(\Gamma)\cup\Gamma)$ $\text{(since}~W~\text{is monotonic)}$
    $=W(\Gamma)$ $\text{(since}~W~\text{satisfies quasi-closure)}$
@@
    $\underline{(3)\implies(2):}$ Suppose that $W$ is not monotonic. So, there exists $\Gamma\cup\Sigma\subseteq\mathscr{L}$ with $\Gamma\subseteq\Sigma$ such that $W(\Gamma)\not\subseteq W(\Sigma)$. This implies that there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W(\Sigma)$. Since $\Gamma\subseteq\Sigma\subseteq W^{\infty}(\Sigma)$ and $\Sigma$ itself is downward $q$-closed by (3), it follows that $W(\Gamma)\subseteq W(\Sigma)$. Hence, $\alpha\in W(\Gamma)$ implies that $\alpha\in W(\Sigma)$ as well - a contradiction. Therefore, $W$ is monotonic.
@@
    $\underline{(4)\implies(2):}$ We first show that $W$ is monotonic. For this, suppose that there exists $\Gamma\cup\Sigma\subseteq\mathscr{L}$ with $\Gamma\subseteq\Sigma$ such that $W(\Gamma)\not\subseteq W(\Sigma)$. This implies that there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W(\Sigma)$. Consequently, by (4), there exists $\Delta\supseteq\Sigma$ such that $\Delta$ is downward $q$-closed and $\alpha\notin W(\Delta)$. Now, since $\Gamma\subseteq\Sigma\subseteq\Delta\subseteq W^{\infty}(\Delta)$, and $\Delta$ is downward $q$-closed, $W(\Gamma)\subseteq W(\Delta)$. So, since $\alpha\in W(\Gamma)$, $\alpha\in W(\Delta)$ as well - a contradiction. Hence, $W$ is monotonic.
    We now show that $W$ satisfies quasi-closure as well. Suppose the contrary, i.e., $W(W^{\infty}(\Gamma))\neq W(\Gamma)$. Then we must have $W\left(W^{\infty}(\Gamma)\right)\not\subseteq W(\Gamma)$ for some $\Gamma\subseteq\mathscr{L}$. So, there exists $\alpha\in W\left(W^{\infty}(\Gamma)\right)$ such that $\alpha\notin W(\Gamma)$. Then, by (4), there exists a downward $q$-closed $\Sigma\supseteq\Gamma$ such that $\alpha\notin W(\Sigma)$. Since $\Sigma$ is downward $q$-closed and $W^{\infty}(\Sigma)\subseteq W^{\infty}(\Sigma)$, it follows that $\alpha\notin W\left(W^{\infty}(\Sigma)\right)$. Now, as $W$ is monotonic, $\Gamma\subseteq\Sigma$ implies that $W^{\infty}(\Gamma)\subseteq W^{\infty}(\Sigma)$, and hence $W\left(W^{\infty}(\Gamma)\right)\subseteq W\left(W^{\infty}(\Sigma)\right)$. Thus, $\alpha\notin W\left(W^{\infty}(\Gamma)\right)$ as well - a contradiction. Hence, $W\left(W^{\infty}(\Gamma)\right)\subseteq W(\Gamma)$. This proves (2).
@@
    The proof that every structural $q$-type logic has a 3-valued description is already been given, e.g., in [30, pp. 204-206]. However, a similar argument works for arbitrary $q$-type logical structures as well. Below we give a proof of the same, to keep the presentation self-contained.